Compute the average atomic mass of gallium, if 60.0% of the gallium atom occurring in nature have mass number 69 and 40.0% of the items have mass number 71.
1
Expert's answer
2020-11-04T14:17:40-0500

Explanation:

Let the percentage of the isotope 71Ga be =x%

Then,

The percentage of the isotope 69Ga is =(100−x)%

We write the mass balance equation

100⋅69.72=x⋅71+(100−x)⋅69

6972=71x+6900−69x

2x=6972−6900=72

x=36

Therefore,

The percentage of the isotope 71Ga is =36%